OpenStudy (anonymous):

There are 6 numbers on a normal dice. What is the probability that if you roll the dice, you'll get a 5 or a 6? Express your answer as a fraction with no spaces. Reduce the fraction to its simplest form.

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110103203708AAe473j View number 9.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4fa00d00e4b029e9dc310a75 This explains why.

OpenStudy (anonymous):

ok and what about this There are 6 numbers on a normal dice. What is the probability that if you roll the dice, you'll get an odd number? Express your answer as a percent. Include the percent sign at the end of your number, and don't put any spaces in your answer.

OpenStudy (anonymous):

3/6 3# =odd 3# =even 3/6=1/2=.5= 50%

OpenStudy (anonymous):

Is that all @adonis201094 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

nvm yes thanks :)

OpenStudy (anonymous):

There are six open containers arranged as shown. A ball is tossed and falls into one of the containers. Find P(a number greater than 4). Write your answer as a fraction (i.e., 1/2).